OpenStudy (anonymous):

What is the midpoint of the segment that has the end points (3a, 5b + 1), (8a-2, 7b)

hartnn (hartnn):

The co-ordinates of midpoint(x,y) on the line segment with end-points (x1,y1) and (x2,y2) is just the average of co-ordinates, given by : \(\large x=\frac{x_1+x_2}{2}\) \(\large y=\frac{y_1+y_2}{2}\) now,just put the values and find the co-ordinates of midpoint (x,y).

OpenStudy (anonymous):

I got 11a-2 divided by two , 12b+1 divided by two but I don't know how to finish dividing it

hartnn (hartnn):

thats your final answer :)

OpenStudy (anonymous):

Oh well thank you

hartnn (hartnn):

welcome ^_^